OpenStudy (s):

Find dy/dx: y=e^(5-3lnx)

OpenStudy (lgbasallote):

use chain rule

OpenStudy (anonymous):

I dont think so he know the chain rule let's help out :)

OpenStudy (s):

so this would be 5/(3lnx)?

OpenStudy (anonymous):

nope answer would be : -3/x(e^5-3lnx)

OpenStudy (s):

could you please show me the steps, i am confused how you get that..

myininaya (myininaya):

@muhe Don't just give the answer. It is against our CoC. Thanks for understanding.

OpenStudy (anonymous):

okay sorry for that I thought he is asking only the answer..

myininaya (myininaya):

\[y=e^{f(x)} => y'=f'(x) e^{f(x)} \]

myininaya (myininaya):

This right here will work every time for something in the form of \[y=e^{f(x)}\]

OpenStudy (anonymous):

U differentiate the power of the e first and then multiply it with the function that is how u differentiate the power of the e..

OpenStudy (s):

sorry, my battery just died

OpenStudy (s):

so, in this case the derivative of 5

myininaya (myininaya):

\[y'=(5-3 \ln(x))'e^{5-3\ln(x)}\] using the rule I wrote down

OpenStudy (s):

the derivative of 5-3lnx would be -3* 1/x ?

myininaya (myininaya):

Yes! So what is y'?

OpenStudy (anonymous):

yes exactly

OpenStudy (s):

so y' would be -3/x * e^(5-3lnx)

myininaya (myininaya):

:) Great job everyone!

OpenStudy (s):

thank you =D now I get that !